A chemist would like to produce a 30% alcohol solution. How many liters of 20% and 50% solution should be mixed in order to get 12 liters of 30% solution?
Answer by robertb(5830)

0.20x + 0.50(12 - x) = 0.30*12
==> 0.20x + 6 - 0.50x = 3.6
==> -0.30x = -2.4
==> x = 8 liters of the 20%, and
12-x = 4 liters of the 50%.